# Truthful Text Sanitization Guided by Inference Attacks

### v1

@@ 3 Text Sanitization Method
Figure 1: Overview of the text sanitization process with INTACT, together with a sample system output.
@@ 3.1 Replacement Generation
In this first step, we prompt an LLM to generate a list of generalized replacement candidates for the provided text span in its context. We explicitly request replacements with different levels of abstraction that the LLM is then required to sort from most to least specific. This sorting is done within the same step in order to limit computational costs. To enhance the quality of the generated candidates, we also include a one-shot example containing a sentence with a target span (of the same label as the span to sanitize) and some possible generalizations. We provide the text span in its local context, defined here as the sentence it occurs in.
More formally, given an open-weights instruction fine-tuned language model $\texttt{LLM}_{\texttt{instr}}$ and an input text span $t_{i}$ occurring in $D$ that should be sanitized, we prompt the LLM to generate a list of $m$ replacements candidates $C_{i}=\{c_{i,1},c_{i,2},\ldots,c_{i,m}\}$, where the candidates $c_{i,j}$ are sorted by order of abstraction (from the most specific to the most generic). The list of replacement candidates $C_{i}$ for the text span $t_{i}$ is thus defined as:
@@
where $p_{C}$ is the prompt used to instruct $\texttt{LLM}_{\texttt{instr}}$ to generate the candidates, $e_{C}$ is the one-shot example, $t_{i}$ is the input text span that should be edited, and $s_{i}$ is the sentence in $D$ containing that text span.
@@
For direct identifiers – such as person names or phone numbers – which can reveal identity even in isolation, replacement candidates are not obtained through Eq. 1. Instead, in line with how direct identifiers are handled in statistical disclosure control (SDC) [19], $C_{i}$ of direct identifiers are assigned to either:
  • 1.
    a generic token (e.g., “***”);
  • 2.
    an entity type (e.g., PERSON) corresponding to the label $l_{i}$;
  • 3.
    a more human-readable variant of this category (e.g., “the person”).
This not only helps maintaining privacy, but it also prevents accidental identification, i.e., generated names or codes matching actual entities.
@@ 3.2 Replacement Selection
The second step of the method identifies the best replacement $c_{i}^{*}$ among the possible replacement candidates $C_{i}$ for the original text span $t_{i}$. The best replacement is here defined as the most specific candidate that does not reveal, directly or indirectly, the original content of the span $t_{i}$. In other words, the best replacement should be the most informative replacement among those that do not constitute a disclosure risk. To determine whether a replacement candidate $c_{i,j}$ constitutes a disclosure risk, we carry out an inference attack – that is, we use the instruction-tuned LLM to produce a set of guesses for the original span given some context. These guesses are then compared to the original span to assess whether the replacement is acceptable (i.e. if none of the guesses match the original span) or whether it would allow an adversary to find back the original text span. Crucially, those guesses are generated by taking into account not only the replacement candidate, but the full (edited) document as context, in order to allow the LLM to take advantage of all information expressed in the document to predict the original content of a given span.
@@
For each text span $t_{i}$ and replacement candidate $c_{i,j}$ for this span, we generate a set of $p$ guesses $G_{i,j}=\{g_{i,j}^{1},g_{i,j}^{2},\allowbreak\ldots,g_{i,j}^{p}\}$ using the instruction-tuned LLM:
@@
    $p_{G}$ is the prompt employed for the generation of guesses.
@@
    $e_{G}$ is an one-shot example composed of a short purpose-written text, with one span substituted with a synthetic LLM replacement and a set of synthetic LLM guesses.
@@
We then need to define what constitutes a “successful” guess from the point of view of the adversary. A naive approach would be to define a guess as successful if it returns the exact same string as the original text span. This approach, however, fails to account for guesses that are closely related but not entirely identical to the original string. To capture those variations, we define a $\small\textsc{match}(t_{i},g_{i,j}^{k})$ condition, which returns true if any of the following conditions hold:
@@
[…]
@@ 3.3 Full sanitization process
The selection of the best replacement $c_{i}^{*}$ among possible alternatives proceeds as follows:
@@
This process is detailed in Algorithm 1.
@@
[…]
@@
     # Generate replacement candidates with prompt $p_{C}$ and one-shot $e_{C}$
@@
[…]
@@
[…]
@@ 4 Evaluation metrics
As discussed in Section 2.3, the current metrics for text anonymization evaluation are not suitable to assess the quality of sanitized replacements. To address this shortcoming, we propose two novel automatic evaluation metrics. Instead of relying on manual annotations, our metrics quantify data utility as a function of the proportion of semantic content preserved in the selected replacements relative to the original text span, and privacy protection as the inverse of the observed residual re-identification risk.
@@ 4.1 Utility assessment
    the information content of the disclosive text span to be protected within the original document, which accounts for the amount of semantics encompassed by the original text span.
@@ 4.1.1 Information content
Various methods are possible to practical compute the probability $\Pr(t)$ in Eq. 3, such as using their frequency of appearance within a corpus [34] (even leveraging the Web as corpus [37, 35]), considering the number of hypernyms and/or hyponyms within an ontology [39, 7], or evaluating their predictability with a language model[32]. The latter approach utilizes the general language understanding capabilities of language models (specifically BERT [12]) to predict words based on context. The model receives the context of the text span (e.g., the surrounding sentence) with the target text span substituted by the [MASK] token. The probability of the text span is then determined by the model’s probability of predicting it from the context. For multi-token text spans, the minimum probability is selected. This predictability technique is the only estimate of $\Pr(t)$ that accounts for compositionality, which states that the meaning of an expression is derived from the meanings of its parts and the rules for combining them. Moreover, this technique is also the only capable of estimating probabilities for any text span, not requiring it to be present in a specific corpus or ontology.
As in [32], we rely on BERT (bert-base-uncased [^4] https://huggingface.co/google-bert/bert-base-uncased) for computing the IC of our text spans. However, contrary to [32], here we must compute the IC for all text spans within a document. Using the method proposed in [32] would involve substituting all text spans with [MASK] simultaneously, depriving the model of the necessary context to make accurate probability predictions. To solve this issue, we alternate the [MASK] substitution across text spans. Specifically, placing [MASK] for one of every $N$ text spans, using the remaining spans as context. This technique requires making multiple predictions, each time shifting the [MASK] placements. For instance, with $N=2$, two predictions are needed: one with [MASK] applied to the odd text spans while using the even spans as context, and another with [MASK] applied to the even spans while using the odd spans as context. Increasing $N$ linearly increments the number of predictions (and thus the computational cost), but it is expected to enhance the quality of the probability estimates by providing more context for each prediction.
@@ 4.1.2 Semantic similarity
As introduced above, the semantic similarity between an original text span and its replacement indicates how well the semantic content of the former has been preserved by the latter. A well-established measure of semantic similarity between text spans consists of calculating the cosine similarity between their embeddings in a language model [43, 11, 4].
A cosine similarity of 1 corresponds to identical spans, while 0 indicates no particular similarity, and -1 depicts completely unrelated text spans. Given that replacements are not expected to be unrelated to the original text span, we constrain the similarity calculation to the $[0,1]$ range:
@@
As embedding model, we leveraged paraphrase-albert-base-v2555https://huggingface.co/sentence-transformers/paraphrase-albert-base-v2 in our experiments, an off-the-shelf language model trained with Sentence-BERT [33] that has been shown to achieve state-of-the-art results in multiple benchmarks for semantic similarity [33, 23].
@@ 4.1.3 Text Preserved Similarity
For unedited text spans, $c_{i}^{*}$ is identical to $t_{i}$, resulting in a similarity of 1 and thus no loss in utility. TPS can be interpreted as the arithmetic mean of semantic similarities for all text spans, weighted by their relative informativeness. The TPS metric range from 0, indicating a loss of all original content, to 1, where $D^{\prime}$ conveys precisely the same meaning as $D$ (e.g., no text spans have been altered).
@@ 4.2 Privacy assessment
The ultimate goal of data anonymization is to minimize the re-identification risk. Considering that the greatest disclosure risk comes from the combination of disclosive text spans [24], a sound privacy evaluation should consider the residual re-identification risk of the chosen replacements as a whole. In [26] we proposed a privacy metric for anonymized text, named text re-identification risk (TRIR) to precisely account for this.
More specifically, TRIR involves simulating a plausible re-identification attack on the protected documents by leveraging, as background knowledge, a collection of public documents of a population of individuals that encompasses the actual individuals to be protected. For example, if the individuals to be protected are hospital patients and the anonymized documents are their medical reports, all the citizens of the hospital’s town could be the population, and the public documents leveraged as background knowledge could be a collection social media posts of those citizens.
To construct the attack, a language model with a classifier head is first trained to exploit this background knowledge to predict the individual corresponding a given document. This classification model is thus trained to identify the information most closely linked to each individual from the public documents. When presented with an anonymized document, the model can then leverage the learned knowledge to re-identify the corresponding individual. If the predicted public individual is in fact the corresponding individual to be protected, re-identification is successful. TRIR is therefore computed as the re-identification accuracy on the set of anonymized documents.
@@ 5.1 Dataset
We employed the documents and manual annotations of the Text Anonymization Benchmark, which encompasses 1,268 English documents extracted from court cases of the European Court of Human Rights (ECHR). The Procedure section of these documents, which is rich in personally identifying information, was manually annotated by 12 law students, who tagged disclosive text spans and specified their corresponding entity type by following privacy-oriented annotation guidelines [32].
On average, 11% of the text spans in the documents were considered as PII by the annotators. The entity types of the corpus are: CODE, (organization), DATETIME, LOC (location), QUANTITY, PERSON, DEM (demographics) and MISC (any personal information not falling in the previous categories). As the proposed approach does not require any fine-tuning, we conduct our experiments on the full set of documents from the TAB corpus. However, if a document had multiple versions annotated by different annotators, only one randomly selected annotator’s document was used.
@@ 5.2 Implementation details
For all semantic categories other than CODE and PERSON, we generate replacement candidates (see Eq. (1)) with a 4-bit quantized Mistral 7B Instruct v.0.2777https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 model with multinomial sampling, a temperature of 0.3 and with a limit of 512 new tokens. The number $m$ of replacements is set to five. The document is segmented into sentences using Spacy [^8] https://spacy.io/. Each target PII span $t_{i}$ is delimited by double square-brackets in the sentence. The complete instruction formatted as a multi-turn chat and the example sentences used are presented in Tables 3 and 2 in the Appendix.
@@
For the CODE and PERSON categories, which are direct identifiers, the replacement consists of the entity type label and a running number (e.g., CODE_0) thereby maintaining reference links between different mentions of the same underlying entity.
We also implement rule-based generalizations for DATE for some common date formats and use LLM-based replacements only as fallback strategy. This is motivated by the well-known weakness of many current smaller-sized LLMs to handle numeric information and the potentially identifying nature of e.g., birth dates. The starting level of abstraction for the list of replacements for dates reflects the specificity of the original date. For example, for a date consisting of day+month+year, the most specific generalizations is month+year, but for an original date span of month+year, the most specific replacement is season+year.
@@
For selecting replacements, we employ the same LLM used for generating the candidates (the aforementioned Mistral 7B instruct model). The number of guesses is set to five and the same delimiters for the span being processed are used as for generation, see the prompt in Table 4 in A.2. When all replacement candidates are guessed, the entity type and a running number (e.g., ORG_2) are used as fallback.
@@
To assess whether the LLM was able to correctly guess the original text span based on its replacement and the rest of the document context, we defined a matching function (see Section 3.2). This function first checks the existence of a lemma overlap (either exact or partial) between the guess and the actual span. Lemma overlap for DATETIME entities is handled somewhat differently, as generalization here is mostly rule-based and results in re-using a subset of the original information (e.g.August 2003 for 3 August 2003). We therefore prompt the LLM to always try to guess exact dates and we only consider a DATETIME span guessed if all lemmas between the original span and a guessed span are the same. Allowing a partial overlap between the original and the guessed lemmas might be an acceptable type and level of generalization also in some other cases involving other entity types (e.g., guessing only the country when the original span specifies both the city and the country). We also check for n-gram overlap if the text span corresponds to a named entity. We choose 4-grams to account for words with the same, longer base (stem), e.g., Turkey - Turkish.

### v2

@@ 3 Text sanitization method
Figure 1: Illustration of text sanitization process with INTACT.
@@ 3.1 Replacement generation
In this first step, we prompt an LLM to generate a list of generalized replacement candidates for the provided text span in its context. We explicitly request replacements with different levels of abstraction that the LLM is then required to sort from most to least specific. This sorting is done within the same step as generation, in order to limit computational costs. To enhance the quality of the generated candidates, we also include a one-shot example containing a sentence with a target span (of the same label as the span to sanitize) and some possible generalizations. We provide the text span in its local context, defined here as the sentence it occurs in. Figure 1 shows an example with (partial) context and sorted replacements for Catholic, where Christian is listed as the most specific replacement and Believer in a particular faith as the least specific one.
Formally, given an instruction fine-tuned language model $\texttt{LLM}_{\texttt{instr}}$ and an input text span $t_{i}$ occurring in $D$ that should be sanitized, we prompt the LLM to generate a list of $m$ replacements candidates $C_{i}=\{c_{i,1},c_{i,2},\ldots,c_{i,m}\}$, where the candidates $c_{i,j}$ are sorted by order of abstraction. Sorting is performed by the LLM during generation, based on the instructions given in the candidate generation prompt (provided in Figure 2), which request that candidates be sorted from most specific to most generic. The list of replacement candidates $C_{i}$ for the text span $t_{i}$ is thus defined as:
@@
where $p_{C}$ is the prompt used to instruct $\texttt{LLM}_{\texttt{instr}}$ to generate the candidates, $e_{C}$ is the one-shot example, $t_{i}$ is the input text span that should be edited, and $s_{i}$ is the sentence in $D$ containing that text span. Including sentence-level context is intended to provide sufficient information to ensure grammatical sanitized outputs and to address potentially ambiguous PII.
@@
For direct identifiers – such as person names or phone numbers – which can reveal identity even in isolation, replacement candidates are not obtained through Eq. 1. Instead, in line with how direct identifiers are handled in statistical disclosure control [30], direct identifiers are replaced by placeholders such as entity labels, for instance PERSON $i$. Those replacements, implemented through heuristic rules, not only help maintaining privacy, but it also prevents unintended identification, i.e.,  names or codes matching actual entities. To preserve co-reference chains, the heuristics assign consistent labels to repeated mentions of the same entity. Rule-based generalisations are also employed to produce consistent generalizations for dates expressed in standard formats, as those can be generated directly from heuristics without relying on LLMs (and thus at a much lower computational cost). See Section 5.2 for more details on those generalization heuristics.
@@ 3.2 Replacement selection
The second step of the method identifies the best replacement $c_{i}^{*}$ among the possible replacement candidates $C_{i}$ for the original text span $t_{i}$. The best replacement is here defined as the most specific candidate that does not reveal, directly or indirectly, the original content of the span $t_{i}$. In other words, the best replacement should be the most informative replacement among those that do not constitute a disclosure risk. To determine whether a replacement candidate $c_{i,j}$ constitutes a disclosure risk, an inference attack is carried out – that is, we use the instruction-tuned LLM to produce a set of guesses for the original span given some context. The inference attacks are performed based on one-shot prompting, without additional fine-tuning of the LLM for this specific task. These guesses are then compared to the original span to assess whether the replacement is acceptable (i.e. if none of the guesses match the original span) or whether it would allow an adversary to find back the original text span. Crucially, those guesses are generated by taking into account not only the replacement candidate, but the full (edited) document as context, in order to allow the LLM to take advantage of all information expressed in the document to predict the original content of a given span. For instance, in the example in Figure 1, the LLM cannot infer Catholic when it is replaced by Monotheist in the original document-long context, it is therefore deemed as a safe replacement.
@@
For each text span $t_{i}$ and replacement candidate $c_{i,j}$ for this span, we generate a set of $p$ guesses $G_{i,j}=\{g_{i,j}^{1},g_{i,j}^{2},\allowbreak\ldots,g_{i,j}^{p}\}$ by prompting the instruction-tuned LLM:
@@
    $p_{G}$ is the one-shot prompt employed for the generation of guesses.
@@
    $e_{G}$ is an one-shot example composed of a short synthetic text, with one span substituted with a synthetic LLM replacement and a set of synthetic LLM guesses.
@@
To evaluate the effectiveness of those attacks, it is necessary to define what constitutes a successful guess from the point of view of the adversary. A naive approach would be to define a guess as successful if it returns the exact same string as the original text span. This approach, however, fails to account for guesses that are closely related but not entirely identical to the original string. To capture those variations, we define a $\small\textsc{match}(t_{i},g_{i,j}^{k})$ condition, which returns true if any of the following conditions hold:
@@
$\small\textsc{risky\_replace}(t_{i},c_{i,j})=\begin{cases}\textsf{true}&\text{if }\exists\ g_{i,j}^{k}\in G_{i,j}\text{ where }\small\textsc{match}(t_{i},g_{i,j}^{k})\!=\!\textsf{true}\\ \textsf{false}&\text{otherwise}\end{cases}$
@@ 3.3 Full sanitization process
The selection of the best replacement $c_{i}^{*}$ among possible alternatives thus proceeds as follows:
@@
This process is detailed in Algorithm 1. For the implementation details, including the prompts used for replacement candidate generation and selection, see Section 5.2 and A.
@@
$T,L\leftarrow\small\textsc{detect\_disclosive\_spans(D)}$
@@
  # Generate sorted candidates with prompt $p_{C}$ and one-shot $e_{C}$
@@
  $c_{i}^{*}\leftarrow\min\left(j\textrm{ with }1\leq j\leq m\text{ and }\small\textsc{risky\_replace}(t_{i},c_{i,j})\!=\!\textsf{false}\right)$
@@
  if $\small\textsc{risky\_replace}(t_{i},c_{i,j})\ \forall\ j:1\leq j\leq m$ then
@@ 4 Evaluation metrics
As discussed in Section 2.3, current evaluation methods for text anonymization mostly focus on the task of detecting the occurrence of PII spans, while the subsequent problem of replacing those spans remains understudied and lacking appropriate evaluation metrics. To address this shortcoming, we propose two novel automatic evaluation metrics, respectively called TPS (Text Preserved Similarity) and TRIR (Text Re-identification Risk). Instead of relying on manual annotations, those metrics quantify data utility as a function of the proportion of semantic content preserved in the selected replacements relative to the original text span, and privacy protection as the inverse of the observed residual re-identification risk.
@@ 4.1 Utility assessment
    the information content of the disclosive text span to be protected within the original document, which accounts for the semantic content encompassed by the original text span.
@@ 4.1.1 Information content
Various methods are possible to practical compute the probability $\Pr(t)$ in Eq. 3, such as using their frequency of appearance within a corpus [49] (even leveraging the Web as corpus [53, 51]), considering the number of hypernyms and/or hyponyms within an ontology [56, 8], or evaluating their predictability with a language model [46]. The latter approach utilizes the general language understanding capabilities of language models (specifically BERT [16]) to predict words based on context. The model receives the context of the text span (e.g., the surrounding sentence) with the target text span substituted by the [MASK] token. The probability of the text span is then determined by the model’s probability of predicting it from the context. This LM-based method makes it possible to estimate the information content of any PII span without the need to rely on an in-domain corpus or ontology. For multi-token text spans, the minimum probability is selected, thereby ensuring that the measure of information content reflects the least predictable (and thus most informative) token.
As in [46], we rely on BERT (bert-base-uncased [^4] https://huggingface.co/google-bert/bert-base-uncased) for computing the IC of our text spans. However, contrary to [46], we must in this case compute the IC for all text spans within a document. Using the method proposed in [46] would involve substituting all text spans with [MASK] simultaneously, depriving the model of the necessary context to make accurate probability predictions. To solve this issue, we alternate the [MASK] substitution across text spans. Specifically, placing [MASK] for one of every $N$ text spans, using the remaining spans as context. This technique requires making multiple predictions, each time shifting the [MASK] placements. For instance, with $N=2$, two predictions are needed: one with [MASK] applied to the odd text spans while using the even spans as context, and another with [MASK] applied to the even spans while using the odd spans as context. Increasing $N$ linearly increments the number of predictions (and thus the computational cost), but also helps refining the probability estimates by providing more context for each prediction.
@@ 4.1.2 Semantic similarity
As mentioned above, the semantic similarity between an original text span and its replacement indicates how well the semantic content of the former is preserved by the latter. A well-established measure of semantic similarity between text spans consists of calculating the cosine similarity between their embeddings derived from a language model [62, 15, 5].
A cosine similarity of 1 corresponds to identical spans, while 0 indicates no particular similarity, and -1 depicts completely unrelated text spans. Given that replacements are related to the original text span, we constrain the similarity calculation to the $[0,1]$ range:
@@
As embedding model, we leverage paraphrase-albert-base-v2555https://huggingface.co/sentence-transformers/paraphrase-albert-base-v2 in our experiments, an off-the-shelf Sentence-BERT model [48] trained in paraphrasing data, which has shown strong results on semantic similarity tasks [48, 35]. Given the proximity between paraphrasing and generalization, this model is well-suited to capture the semantic shifts resulting from text sanitization.
@@ 4.1.3 Text preserved similarity
For unedited text spans, $c_{i}^{*}$ is identical to $t_{i}$, resulting in a similarity of 1 and thus no loss in utility. TPS can be interpreted as the arithmetic mean of semantic similarities for all text spans, weighted by their relative informativeness. The TPS metric ranges from 0, indicating a loss of all original content, to 1, where $D^{\prime}$ conveys precisely the same meaning as $D$ (e.g., no text spans have been altered).
@@ 4.2 Privacy assessment
The ultimate goal of data anonymization is to minimize the re-identification risk. Considering that the greatest disclosure risk comes from the combination of disclosive text spans [36], a sound privacy evaluation should consider the residual re-identification risk of the chosen replacements as a whole. In structured data, this is done through record linkage attacks, which are the standard for empirical risk assessment [1, 17]. Building on this concept, [38] introduced an automatic metric called text re-identification risk (TRIR), which adapts the principles of record linkage to text, providing an objective way to assess the residual re-identification risk of anonymized documents.
More specifically, TRIR involves simulating a re-identification attack on the protected documents by leveraging, as background knowledge, a comprehensive collection of public documents of a population of individuals that encompasses the actual individuals to be protected. For example, if the individuals to be protected are hospital patients and the anonymized documents are their medical reports, all the citizens of the hospital’s town could be the population, and the public documents leveraged as background knowledge could be a collection social media posts of those citizens.
To construct the attack, a language model with a classifier head is first trained to exploit this background knowledge to predict the individual corresponding a given document. This classification model is thus trained to identify the information most closely linked to each individual from the public documents. When presented with an anonymized document, the model can then leverage the learned knowledge to re-identify the corresponding individual. If the predicted public individual is in fact the corresponding individual to be protected, re-identification is successful. TRIR is computed as the re-identification accuracy on the set of anonymized documents, which is in line to how record linkage is quantified for tabular data. Importantly, TRIR does not assume that the protected document is an abridged version of an original document, but any text containing information about an individual. This ensures full compatibility with text sanitization strategies..
@@ 5.1 Dataset
We employed the documents and manual annotations of the Text Anonymization Benchmark, which encompasses 1,268 English documents extracted from court cases of the European Court of Human Rights (ECHR). The Procedure section of these documents, which is particularly rich in PII of various types, was manually annotated by 12 law students, who tagged disclosive text spans and specified their corresponding entity type by following privacy-oriented annotation guidelines [46].
On average, 11% of the text spans in the documents were considered as PII by the annotators. The entity types of the corpus are: CODE, ORG (organization), DATETIME, LOC (location), QUANTITY, PERSON, DEM (demographics) and MISC (any personal information not falling in the previous categories). As the proposed approach does not require any fine-tuning, we conduct our experiments on the full set of documents from the TAB corpus. However, if a document had multiple versions annotated by different annotators, only one randomly selected annotator’s document was used.
@@ 5.2 Implementation details
For all semantic categories other than CODE and PERSON, we generate replacement candidates (see Eq. (1)) with a 4-bit quantized Mistral 7B Instruct v.0.2777https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 model with multinomial sampling, a temperature of 0.3 and with a limit of 512 new tokens. The number $m$ of replacements is set to five. The document is segmented into sentences using spaCy [^8] https://spacy.io/. Each target PII span $t_{i}$ is delimited by double square-brackets in the sentence. The prompt used for generating replacement candidates in a one-shot setting is presented in Figure 2. It is set up as a multi-turn chat to accommodate the example containing a matching entity type compared to the target span (the examples are listed in Table 4 in the Appendix). Given that prompt variations considerably influence LLM outputs [47], we considered a number of prompt alternatives. Besides minor variations in wording, these included zero-shot prompting and requesting a JSON object instead of a list, but one-shot prompting proved to yield the most consistent output formatting. Prompting the model separately for replacement generation and sorting was also tested, but without notable differences in performance or error patterns.
Figure 2: One-shot prompt for replacement generation.
Role Message
Here is an example:
{Example sentence context}
Sorted replacements for {[[Example target span]]}:
Assistant {Example of sorted replacements}
User Now do the same for the sentence below.
Original:
{Sentence context}
Sorted replacements for {[[Target span]]}:
@@
For the CODE and PERSON categories, which are direct identifiers, the replacement consists of the entity type label and a running number (e.g., CODE_1) thereby maintaining reference links between different mentions of the same underlying entity. [^9] While direct identifiers could alternatively be replaced by terms including quasi-identifiers already present in the text (e.g. “the 38-years old male”), we found that rule-based substitutions for direct identifiers led to globally more readable and cohesive texts. We also implement rule-based generalizations for DATE for standard date formats. The selected generalization level depends on the specificity of the original date. For example, for a date consisting of day+month+year, the most specific generalizations is month+year, but for an original date span of month+year, the most specific replacement is season+year. Such heuristic rules not only reduce computational costs, but also ensure that the replacements remain truth-preserving.
@@
For selecting replacements, we employ the same LLM used for generating the candidates (the aforementioned Mistral 7B instruct model) without additional fine-tuning. The number of guesses is set to five and the same delimiters for the span being processed are used as for generation. Figure 3 provides the one-shot prompt used for the selection of replacement candidates based on inference attacks. Here, the same example was used in all cases. When all replacement candidates are guessed, the entity type and a running number (e.g., ORG_2) are used as fallback.
Figure 3: Prompt for conducting inference attacks (i.e. guessing the original PII span based on a replacement candidate and the rest of the document).
Role Message
Text: “early 1980s” -$\rangle$ Guess: “3 September 1981”. List more likely guesses first. Do not return anything else. See the example below.
Guesses for “a European Research Institute”:
- the Institute of Polish Literature, University of Warsaw
- the Center for European Neighborhood Studies
- the Centre for Research on the History and Culture of Eastern Europe
- the Leibniz Institute for the History and Culture of Eastern Europe
User Do the same for the text below.
Text: {Text with replacements}
Guesses for {[[Target span]]}:
@@
To assess whether the LLM was able to correctly guess the original text span based on its replacement and the rest of the document context, we rely on the matching function described in Section 3.2. DATETIME entities are handled somewhat differently, as the rule-based generalization of those entities often leads to replacements that include a subset of the lemmas from the original span (e.g.August 2003 for 3 August 2003). We therefore prompt the LLM to always guess exact dates and we only consider a DATETIME span guessed if all lemmas between the original span and a guessed span are the same. Allowing a partial overlap between the original and the guessed lemmas might be an acceptable type and level of generalization also in some other cases involving other entity types (e.g., guessing only the country when the original span specifies both the city and the country). We also check for n-gram overlap if the text span corresponds to a named entity. We choose 4-grams to account for words with the same, longer base (stem), e.g., Turkey - Turkish.